An object is launched straight up into the air at an initial velocity of 64 feet per second. It is launched from a height of 6 feet off the ground. Its height H, in feet, at t seconds is given by the equation H = -16t^2 + 64t + 6. Find all times t that the object is at a height of 54 feet off the ground.
Answer by Alan3354(69443)

H = -16t^2 + 64t + 6 = 54
-t^2 + 4t - 3 = 0
-(t-1)*(t-3) = 0
t = 1 second ascending
t = 3 seconds descending